The product of a number and 9, less 20 is 7.
What is that asking, and what is the answer?

Found 2 solutions by rfer, Theo:
Answer by rfer(16322)

(9x)-20=7
9x=7+20
x=27/9
x=3

Answer by Theo(13342)

they are saying that The product of a number and 9, less 20 is 7.

they want you to solve for the number.

let x = the number.

that statement can be trsnslated algebraically to the following:

9 * x - 20 = 7

add 20 to both sides of that equation to get 9 * x = 27

divide both sides of that equation by 9 to get x = 27/9 = 3

the number is 3.

go back to the original statement and see if it makes sense when the number is 3.

the original statement says:

The product of a number and 9, less 20 is 7.

the product of the number and 9 is 3 * 9 = 27

less 20 = 7

statement makes sense and is true when the number is equal to 3.